# Central limit theorems for interacting innovation processes, related statistical tools and general results

### v3

@@ 2. Central limit theorems
where $\boldsymbol{1}$}} denotes the vector with all the components equal to $1$. As observed above, the matrix $\Gamma$ rules the production of potential novelties and, in particular, its elements out of the diagonal regulate the interaction among the innovation processes with respect to this issue; while, the matrix $W$ rules the interaction among the innovation processes with respect to the choice of an old item. Moreover, set ${\mathbf{D}}^{*}_{t}=(D^{*}_{t,1},\dots,D^{*}_{t,N})^{\top}$, ${\mathbf{D}}^{**}_{\infty}=(D^{**}_{\infty,1},\dots,D^{**}_{\infty,N})^{\top}$, ${\mathbf{P}}_{t}(c)=(P_{t}(1,c),\dots,P_{t}(N,c))^{\top}$ and finally ${\mathbf{K}}_{t}(c)=(K_{t}(1,c),\dots,K_{t}(N,c))^{\top}$.
@@
Assume that the non-negative matrix $\Gamma=(\gamma_{j,h})_{j,h=1,\dots,N}$ is irreducible. Denote by $\gamma^{*}\in(0,1)$ the Perron-Frobenius eigenvalue of $\Gamma$, by $\mathbf{v}$}} the corresponding right eigenvector with strictly positive entries and such that ${{\mathbf{v}}}^{\top}{{\boldsymbol{1}}}=1$ and denote by $\mathbf{u}$}} the corresponding left eigenvector with strictly positive entries and such that ${{\mathbf{v}}}^{\top}{{\mathbf{u}}}=1$. Finally, let $\gamma^{*}_{2}$ be an eigenvalue of $\Gamma$ different from $\gamma^{*}$ and with highest real part and assume that ${\mathcal{R}e}(\gamma^{*}_{2})/\gamma^{*}<1/2$.
@@
Assume that the non-negative matrix $W=(w_{j,h})_{j,h=1,\dots,N}$ is irreducible. Denote by ${\mathbf{v}}_{W}$} the right eigenvector associated to the leading eigenvalue $w^{*}=1$, with strictly positive entries and such that ${{\mathbf{v}}_{W}}^{\top}{{\boldsymbol{1}}}=1$. Finally, let $w^{*}_{2}$ be an eigenvalue of $W$ different from $w^{*}=1$ and with highest real part and assume that ${\mathcal{R}e}(w^{*}_{2})<\frac{1}{2}$.
We are now ready to provide the central limit theorems for the considered model, that describe the second-order behaviour of the quantities in Theorems 1.1 and 1.2. The symbol $\boldsymbol{0}$}} denotes the vector with all the components equal to 0.
@@
$t^{\gamma^{*}/2}\left(\frac{{\mathbf{D}}_{t}^{*}}{t^{\gamma^{*}}}-{\mathbf{D}}^{**}_{\infty}\right)\stackrel{{\scriptstyle stably}}{{\longrightarrow}}\mathcal{N}({\boldsymbol{0}},\widetilde{Z}^{**}_{\infty}C_{det,\Gamma})\,$,
where $\widetilde{Z}_{\infty}^{**}$ is an integrable strictly positive random variable such that ${\mathbf{D}}^{**}_{\infty}=\frac{\widetilde{Z}^{**}_{\infty}}{\gamma^{*}}\,{\mathbf{u}}$ and $C_{det,\Gamma}$ is a deterministic matrix depending on the eigen-structure of $\Gamma$, which is possible to write explicitly.
@@
where $r_{t,h}=1/(\theta_{h}+t+1)=1/(t+1)+O(1/t^{2})$  [^2] In all the paper, the notation $O(s_{t})$ denotes a generic (possibly random) remainder term $R_{t}$ such that $|R_{t}|\leq Cs_{t}$ for a suitable constant $C$ and for $t$ large enough.. The corresponding vectorial dynamics for ${\mathbf{Z}}^{*}_{t}=(Z^{*}_{t,1},\dots,Z^{*}_{t,N})^{\top}$ is
@@
More precisely, from (Gouet, 1993, Lemma 4.1) we have
@@
$\begin{split}\frac{{\mathbf{D}}^{*}_{0}}{\zeta_{0}(\gamma^{*})}&={\boldsymbol{0}}\\ \frac{{\mathbf{D}}^{*}_{t+1}}{\zeta_{t+1}(\gamma^{*})}&=\frac{\zeta_{t}(\gamma^{*})}{\zeta_{t+1}(\gamma^{*})}\frac{{\mathbf{D}}_{t}^{*}}{\zeta_{t}(\gamma^{*})}+\frac{1}{\zeta_{t+1}(\gamma^{*})}{\mathbf{X}}^{*}_{t+1}\\ &=\left(1-\frac{\gamma^{*}}{t+1}\right)\frac{{\mathbf{D}}_{t}^{*}}{\zeta_{t}(\gamma^{*})}+\frac{1}{\zeta_{t+1}(\gamma^{*})}\Delta{\mathbf{M}}^{*}_{t+1}+\frac{1}{\zeta_{t+1}(\gamma^{*})}{\mathbf{Z}}^{*}_{t}\\ &=\left(1-\frac{\gamma^{*}}{t+1}\right)\frac{{\mathbf{D}}_{t}^{*}}{\zeta_{t}(\gamma^{*})}+\frac{1}{\zeta_{t+1}(\gamma^{*})\zeta_{t+1}(1-\gamma^{*})}\Delta{\mathbf{M}}^{**}_{t+1}\\ &+\frac{1}{\zeta_{t+1}(\gamma^{*})\zeta_{t+1}(1-\gamma^{*})}\frac{\zeta_{t+1}(1-\gamma^{*})}{\zeta_{t}(1-\gamma^{*})}{\mathbf{Z}}^{**}_{t}\,.\end{split}$
@@
where $O(\zeta_{t}(1-\gamma^{*})/t^{2})=O(1/t^{1+\gamma^{*}})$. Moreover, we recall that in (Aletti et al., 2023, Theorem S1.1 and Lemma S1.2) we proved
@@
where $\widetilde{Z}_{\infty}^{**}$ is an integrable strictly positive random variable and ${\mathbf{D}}^{**}_{\infty}=\frac{\widetilde{Z}^{**}_{\infty}}{\gamma^{*}}\,{\mathbf{u}}$. Therefore, we can apply Theorem A.4 with ${\mathbf{A}}_{t}={\mathbf{Z}}^{**}_{t}=\zeta_{t}(1-\gamma^{*}){\mathbf{Z}}^{*}_{t}$, ${\mathbf{B}}_{t}=\frac{{\mathbf{D}}_{t}^{*}}{\zeta_{t}(\gamma^{*})}$, $\Phi=\Gamma$, $\phi^{*}=\gamma^{*}$, $\Sigma_{\infty}=\widetilde{Z}^{**}_{\infty}diag({\mathbf{u}})$, $\Delta{{\mathbf{M}}}_{t+1}=\Delta{\mathbf{M}}^{**}_{t+1}=\zeta_{t}(1-\gamma^{*})\Delta{\mathbf{M}}^{*}_{t+1}=\zeta_{t}(1-\gamma^{*})({\mathbf{X}}^{*}_{t+1}-{\mathbf{Z}}^{*}_{t})$. Finally, we can conclude by observing that, since $\gamma^{*}<1$, we have
@@
$t^{\gamma^{*}/2}\left(\frac{{\mathbf{D}}_{t}^{*}}{\sum_{h=1}^{N}D^{*}_{t,h}}-\frac{{\mathbf{u}}}{\sum_{h=1}^{N}u_{h}}\right)\stackrel{{\scriptstyle stably}}{{\longrightarrow}}\mathcal{N}\left({\boldsymbol{0}},\frac{1}{\widetilde{Z}^{**}_{\infty}}\frac{(\gamma^{*})^{2}}{(\sum_{j=1}^{N}u_{j})^{2}}Q_{\Gamma}C_{det,\Gamma}Q_{\Gamma}^{\top}\right)\,$,
@@
$t^{\gamma^{*}/2}\left(\frac{{\mathbf{D}}_{t}^{*}}{D^{*}_{t,j}}-\frac{{\mathbf{u}}}{u_{j}}\right)\stackrel{{\scriptstyle stably}}{{\longrightarrow}}\mathcal{N}\left({\boldsymbol{0}},\frac{1}{\widetilde{Z}^{**}_{\infty}}\frac{(\gamma^{*})^{2}}{u_{j}^{2}}Q_{j}C_{det,\Gamma}Q_{j}^{\top}\right)\,$,
@@
We have to apply Corollary A.6 reported in the appendix with ${\mathbf{A}}_{t}={\mathbf{Z}}^{**}_{t}=\zeta_{t}(1-\gamma^{*}){\mathbf{Z}}^{*}_{t}$, ${\mathbf{B}}_{t}=\frac{{\mathbf{D}}_{t}^{*}}{\zeta_{t}(\gamma^{*})}$, ${\mathbf{x}}={\boldsymbol{1}}$, $\widetilde{A}_{\infty}=\widetilde{Z}_{\infty}^{**}$ and $\phi^{*}=\gamma^{*}$. Note that, as specified in the proof of this corollary, we have $\widetilde{Z}_{\infty}^{**}Q_{\Gamma}C_{det,\Gamma}Q_{\Gamma}^{\top}=Q_{\Gamma}{\mathcal{M}}^{33}_{\infty}Q_{\Gamma}^{\top}$, where ${\mathcal{M}}^{33}_{\infty}$ is the matrix defined in Theorem A.3] when $N\geq 2$ with the matrix $\Phi$ equal to $\Gamma$. Similarly, we obtain the second part of the statement by means of Corollary A.6 with ${\mathbf{x}}={\mathbf{e}}_{j}$. ∎
@@
Under Assumption 2.2, we have for each observed color $c$ in the system
@@
For any color $c$ already observed in the system at time $t$, we can write
@@
for $t\geq t^{*}(c)$, where $t^{*}(c)$ is the time-step of the first extraction of $c$ in the system, ${\Delta{\mathbf{K}}}_{t}(c)=(\Delta K_{t}(1,c),\dots,\Delta K_{t}(N,c))^{\top}$ and $\Delta{{\mathbf{M}}}_{t+1}(c)={\Delta{\mathbf{K}}}_{t+1}(c)-{\mathbf{P}}_{t}(c)$.
Considering also the dynamics of ${\mathbf{K}}_{t}(c)/t$, we obtain for $t\geq t^{*}(c)$
@@
Moreover, from Theorem 1.2, we have $\widetilde{P}_{t}(c)={\mathbf{v}}_{W}^{\top}{\mathbf{P}}_{t}(c)\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\widetilde{P}_{\infty}(c)\,$, and this convergence is also in mean because $(\widetilde{P}_{t}(c))_{t}$ is uniformly bounded. Therefore, we can apply Theorem A.4 with ${\mathbf{A}}_{t}={\mathbf{P}}_{t}(c)$, ${\mathbf{B}}_{t}={\mathbf{K}}_{t}(c)/t$, $\Phi=W$, $\phi^{*}=w^{*}=1$, ${\mathbf{u}}={\mathbf{u}}_{W}={\boldsymbol{1}}$, ${\mathbf{v}}={\mathbf{v}}_{W}$, $\Sigma_{\infty}=\widetilde{P}_{\infty}(c)(1-\widetilde{P}_{\infty}(c))I$, $\Delta{{\mathbf{M}}}_{t+1}=\Delta{{\mathbf{M}}}_{t+1}(c)={\Delta{\mathbf{K}}}_{t+1}(c)-{\mathbf{P}}_{t}(c)$. Note that in Theorem A.4 we have the explicit formulas for the elements of the matrix $C_{det,W}$: that is, $C_{det,W}=1$ when $N=1$ and
@@
Under Assumption 2.2, we have for each observed color $c$ in the system
@@
When ${\mathcal{R}e}(\gamma^{*}_{2})/\gamma^{*}\geq 1/2$, as highlighted in Remark A.1, the rate of convergence is slower than $t^{\gamma^{*}/2}$. Indeed, similarly to the stochastic approximation literature (e.g. Zhang (2016)), although we deal with a random limit covariance matrix, for the case ${\mathcal{R}e}(\gamma^{*}_{2})/\gamma^{*}=1/2$ we conjecture a CLT with a logarithmic correction in the scaling; while in the case ${\mathcal{R}e}(\gamma^{*}_{2})/\gamma^{*}>1/2$ we expect a power-law rate with an exponent smaller than $\gamma^{*}/2$, but not a Gaussian limit. Analogous considerations can be done for the cases ${\mathcal{R}e}(w^{*}_{2})=1/2$ and ${\mathcal{R}e}(w^{*}_{2})>1/2$.

### v4

@@ 3. Central limit theorems
where $\boldsymbol{1}$}} denotes the vector with all the components equal to $1$. As observed above, the matrix $\Gamma$ rules the production of potential novelties and, in particular, its elements out of the diagonal regulate the interaction among the innovation processes with respect to this issue; while, the matrix $W$ rules the interaction among the innovation processes with respect to the choice of an old item. Moreover, recall that ${\mathbf{D}}^{*}_{t}=(D^{*}_{t,1},\dots,D^{*}_{t,N})^{\top}$, ${\mathbf{P}}_{t}(c)=(P_{t}(1,c),\dots,P_{t}(N,c))^{\top}$ and ${\mathbf{K}}_{t}(c)=(K_{t}(1,c),\dots,K_{t}(N,c))^{\top}$. Finally, the symbol $\boldsymbol{0}$}} will denote the vector with all the components equal to 0.
@@
Assume that the non-negative matrix $\Gamma=(\gamma_{j,h})_{j,h=1,\dots,N}$ is irreducible and diagonalizable. Denote by $\gamma^{*}\in(0,1)$ the Perron-Frobenius eigenvalue of $\Gamma$, by $\mathbf{v}$}} the corresponding right eigenvector with strictly positive entries and such that ${{\mathbf{v}}}^{\top}{{\boldsymbol{1}}}=1$ and denote by $\mathbf{u}$}} the corresponding left eigenvector with strictly positive entries and such that ${{\mathbf{v}}}^{\top}{{\mathbf{u}}}=1$. Finally, let $\gamma^{*}_{2}$ be an eigenvalue of $\Gamma$ different from $\gamma^{*}$ and with highest real part and assume that ${\mathcal{R}e}(\gamma^{*}_{2})/\gamma^{*}<1/2$.
@@
Assume that the non-negative matrix $W=(w_{j,h})_{j,h=1,\dots,N}$ is irreducible and diagonalizable. Denote by ${\mathbf{v}}_{W}$} the right eigenvector associated to the leading eigenvalue $w^{*}=1$, with strictly positive entries and such that ${{\mathbf{v}}_{W}}^{\top}{{\boldsymbol{1}}}=1$. Finally, let $w^{*}_{2}$ be an eigenvalue of $W$ different from $w^{*}=1$ and with highest real part and assume that ${\mathcal{R}e}(w^{*}_{2})<\frac{1}{2}$.
We are now ready to provide the central limit theorems for the considered model, that describe the second-order behaviour of the quantities in Theorems 2.1 and 2.2 under the above assumptions. A discussion of possible extensions of these results is postponed to Remark 3.1, after the proofs, at the end of this section. For completeness, we also point out that the special case $N=1$ is investigated in Bercu and Favaro, (2024); Gouet, (1993).
@@
$t^{\gamma^{*}/2}\left(\frac{{\mathbf{D}}_{t}^{*}}{t^{\gamma^{*}}}-D^{**}_{\infty}{\mathbf{u}}\right)\stackrel{{\scriptstyle stably}}{{\longrightarrow}}\mathcal{N}({\boldsymbol{0}},D^{**}_{\infty}\gamma^{*}C_{det,\Gamma})\,$,
where $D^{**}_{\infty}$ is defined in Theorem 2.1 and $C_{det,\Gamma}$ is a deterministic matrix depending on the eigen-structure of $\Gamma$, which is possible to write explicitly.
@@
where $r_{t,h}=1/(\theta_{h}+t+1)=1/(t+1)+O(1/t^{2})$  [^2] In all the paper, the notation $O(s_{t})$ denotes a generic (possibly random) remainder term $R_{t}$ such that $|R_{t}|\leq Cs_{t}$ for a suitable deterministic constant $C$ and for $t$ large enough.. The corresponding vectorial dynamics for ${\mathbf{Z}}^{*}_{t}=(Z^{*}_{t,1},\dots,Z^{*}_{t,N})^{\top}$ is
@@
More precisely, from (Gouet,, 1993, Lemma 4.1) we have
@@
$\begin{split}\frac{{\mathbf{D}}^{*}_{0}}{\zeta_{0}(\gamma^{*})}&={\boldsymbol{0}}\\ \frac{{\mathbf{D}}^{*}_{t+1}}{\zeta_{t+1}(\gamma^{*})}&=\frac{\zeta_{t}(\gamma^{*})}{\zeta_{t+1}(\gamma^{*})}\frac{{\mathbf{D}}_{t}^{*}}{\zeta_{t}(\gamma^{*})}+\frac{1}{\zeta_{t+1}(\gamma^{*})}{\mathbf{X}}^{*}_{t+1}\\ &=\left(1-\frac{\gamma^{*}}{t+1}+O(t^{-2})\right)\frac{{\mathbf{D}}_{t}^{*}}{\zeta_{t}(\gamma^{*})}+\frac{1}{\zeta_{t+1}(\gamma^{*})}\Delta{\mathbf{M}}^{*}_{t+1}+\frac{1}{\zeta_{t+1}(\gamma^{*})}{\mathbf{Z}}^{*}_{t}\\ &=\left(1-\frac{\gamma^{*}}{t+1}+O(t^{-2})\right)\frac{{\mathbf{D}}_{t}^{*}}{\zeta_{t}(\gamma^{*})}+\frac{1}{\zeta_{t+1}(\gamma^{*})\zeta_{t+1}(1-\gamma^{*})}\Delta{\mathbf{M}}^{**}_{t+1}\\ &+\frac{1}{\zeta_{t+1}(\gamma^{*})\zeta_{t+1}(1-\gamma^{*})}\frac{\zeta_{t+1}(1-\gamma^{*})}{\zeta_{t}(1-\gamma^{*})}{\mathbf{Z}}^{**}_{t}\,.\end{split}$
@@
where $O(\zeta_{t}(1-\gamma^{*})/t^{2})=O(1/t^{1+\gamma^{*}})$. Moreover, we recall that in (Aletti et al.,, 2023, Theorem S1.1 and Lemma S1.2) (see also [8, Appendix B.1] for a general theory) we proved
@@
where $\widetilde{Z}_{\infty}^{**}$ is an integrable strictly positive random variable and ${\mathbf{Z}}_{t}^{**}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\widetilde{Z}_{\infty}^{**}{\mathbf{u}}$ so that $D_{\infty}^{**}=\widetilde{Z}^{**}_{\infty}/\gamma^{*}$. In addition, we have
$t^{-(1-\gamma^{*})}E[\Delta{\mathbf{M}}^{**}_{t+1}(\Delta{\mathbf{M}}^{**}_{t+1})^{\top}\mid\mathcal{F}_{t}]$ $=\frac{\zeta_{t}(1-\gamma^{*})^{2}}{t^{1-\gamma^{*}}}E[\Delta{\mathbf{M}}^{*}_{t+1}(\Delta{\mathbf{M}}^{*}_{t+1})^{\top}\mid\mathcal{F}_{t}]$
$=\frac{\zeta_{t}(1-\gamma^{*})^{2}}{t^{1-\gamma^{*}}}diag({\mathbf{Z}}^{*}_{t})(I-diag({\mathbf{Z}}^{*}_{t}))$
$=\frac{\zeta_{t}(1-\gamma^{*})}{t^{1-\gamma^{*}}}diag({\mathbf{Z}}^{**}_{t})(I-diag({\mathbf{Z}}^{*}_{t}))\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\Sigma_{\infty}=\widetilde{Z}^{**}_{\infty}diag({\mathbf{u}}).$
Therefore, we can apply Theorem A.4 with ${\mathbf{A}}_{t}={\mathbf{Z}}^{**}_{t}=\zeta_{t}(1-\gamma^{*}){\mathbf{Z}}^{*}_{t}$, ${\mathbf{B}}_{t}=\frac{{\mathbf{D}}_{t}^{*}}{\zeta_{t}(\gamma^{*})}$, $\Phi=\Gamma$, $\phi^{*}=\gamma^{*}$, $\Sigma_{\infty}=\widetilde{Z}^{**}_{\infty}diag({\mathbf{u}})$, $\Delta{{\mathbf{M}}}_{t+1}=\Delta{\mathbf{M}}^{**}_{t+1}=\zeta_{t}(1-\gamma^{*})\Delta{\mathbf{M}}^{*}_{t+1}=\zeta_{t}(1-\gamma^{*})({\mathbf{X}}^{*}_{t+1}-{\mathbf{Z}}^{*}_{t})$. Finally, we can conclude by observing that, since $\gamma^{*}<1$, we have
@@
$t^{\gamma^{*}/2}\left(\frac{{\mathbf{D}}_{t}^{*}}{\sum_{h=1}^{N}D^{*}_{t,h}}-\frac{{\mathbf{u}}}{\sum_{h=1}^{N}u_{h}}\right)\stackrel{{\scriptstyle stably}}{{\longrightarrow}}\mathcal{N}\left({\boldsymbol{0}},\frac{1}{D^{**}_{\infty}}\frac{\gamma^{*}}{(\sum_{j=1}^{N}u_{j})^{2}}Q_{\Gamma}C_{det,\Gamma}Q_{\Gamma}^{\top}\right)\,$,
@@
$t^{\gamma^{*}/2}\left(\frac{{\mathbf{D}}_{t}^{*}}{D^{*}_{t,j}}-\frac{{\mathbf{u}}}{u_{j}}\right)\stackrel{{\scriptstyle stably}}{{\longrightarrow}}\mathcal{N}\left({\boldsymbol{0}},\frac{1}{D^{**}_{\infty}}\frac{\gamma^{*}}{u_{j}^{2}}Q_{j}C_{det,\Gamma}Q_{j}^{\top}\right)\,$,
@@
We have to apply Corollary A.6 reported in the appendix with ${\mathbf{A}}_{t}={\mathbf{Z}}^{**}_{t}=\zeta_{t}(1-\gamma^{*}){\mathbf{Z}}^{*}_{t}$, ${\mathbf{B}}_{t}=\frac{{\mathbf{D}}_{t}^{*}}{\zeta_{t}(\gamma^{*})}$, ${\mathbf{x}}={\boldsymbol{1}}$, $\widetilde{A}_{\infty}=\widetilde{Z}_{\infty}^{**}=\gamma^{*}D^{**}_{\infty}$ and $\phi^{*}=\gamma^{*}$. Note that, as specified in the proof of this corollary, we have $\widetilde{Z}_{\infty}^{**}Q_{\Gamma}C_{det,\Gamma}Q_{\Gamma}^{\top}=Q_{\Gamma}{\mathcal{M}}^{33}_{\infty}Q_{\Gamma}^{\top}$, where ${\mathcal{M}}^{33}_{\infty}$ is the matrix defined in Theorem A.3] when $N\geq 2$ with the matrix $\Phi$ equal to $\Gamma$. Similarly, we obtain the second part of the statement by means of Corollary A.6 with ${\mathbf{x}}={\mathbf{e}}_{j}$. ∎
@@
Under Assumption 3.2, we have for a given observed color $c$ in the system
@@
For a given color $c$, let $t^{*}(c)$ be the random time-step of the first extraction of $c$ in the system. Conditionally on $\{t^{*}(c)<+\infty\}$, for $t\geq t^{*}(c)$ we can write
@@
where ${\Delta{\mathbf{K}}}_{t}(c)=(\Delta K_{t}(1,c),\dots,\Delta K_{t}(N,c))^{\top}$ and $\Delta{{\mathbf{M}}}_{t+1}(c)={\Delta{\mathbf{K}}}_{t+1}(c)-{\mathbf{P}}_{t}(c)$.
Considering also the dynamics of ${\mathbf{K}}_{t}(c)/t$, we obtain
@@
Moreover, from Theorem 2.2, we have $\widetilde{P}_{t}(c)={\mathbf{v}}_{W}^{\top}{\mathbf{P}}_{t}(c)\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\widetilde{P}_{\infty}(c)\,$, and this convergence is also in mean because $(\widetilde{P}_{t}(c))_{t}$ is uniformly bounded. In addition, we have
$E[\Delta{\mathbf{M}}_{t+1}(\Delta{\mathbf{M}}_{t+1})^{\top}\mid\mathcal{F}_{t}]=diag({\mathbf{P}}_{t})(I-diag({\mathbf{P}}_{t}))\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\Sigma_{\infty}=\widetilde{P}_{\infty}(c)(1-\widetilde{P}_{\infty}(c))I.$
Therefore, we can apply Theorem A.4 with ${\mathbf{A}}_{t}={\mathbf{P}}_{t}(c)$, ${\mathbf{B}}_{t}={\mathbf{K}}_{t}(c)/t$, $\Phi=W$, $\phi^{*}=w^{*}=1$, ${\mathbf{u}}={\mathbf{u}}_{W}={\boldsymbol{1}}$, ${\mathbf{v}}={\mathbf{v}}_{W}$, $\Sigma_{\infty}=\widetilde{P}_{\infty}(c)(1-\widetilde{P}_{\infty}(c))I$, $\Delta{{\mathbf{M}}}_{t+1}=\Delta{{\mathbf{M}}}_{t+1}(c)={\Delta{\mathbf{K}}}_{t+1}(c)-{\mathbf{P}}_{t}(c)$. Note that in Theorem A.4 we have the explicit formulas for the elements of the matrix $C_{det,W}$: that is, $C_{det,W}=1$ when $N=1$ and
@@
Under Assumption 3.2, we have for a given observed color $c$ in the system
@@
(Discussion of the Assumption 3.1 and Assumption 3.2)
As observed in Remark A.1 in the appendix, removing the diagonalizability of the matrices $\Gamma$ and $W$ leads to analogous results (cf. Yang et al., (2025)). Moreover, in order to deal with reducible matrices, we refer to Aletti et al., (2023, 2024); Aletti and Ghiglietti, (2017). Finally, when ${\mathcal{R}e}(\gamma^{*}_{2})/\gamma^{*}\geq 1/2$, as highlighted in Remark A.2 in appendix, the rate of convergence is slower than $t^{\gamma^{*}/2}$. Indeed, similarly to the stochastic approximation literature (e.g. Zhang, (2016)), for the case ${\mathcal{R}e}(\gamma^{*}_{2})/\gamma^{*}=1/2$ we expect a CLT with a logarithmic correction in the scaling; while in the case ${\mathcal{R}e}(\gamma^{*}_{2})/\gamma^{*}>1/2$, a power-law rate with an exponent smaller than $\gamma^{*}/2$ could hold true, but not a Gaussian limit. Analogous considerations can be done for the cases ${\mathcal{R}e}(w^{*}_{2})=1/2$ and ${\mathcal{R}e}(w^{*}_{2})>1/2$.